A 33.5 Ω resistor is connected to a power supply. As a result there is a current of 12.2 mA flowing through the resistor. Determine the power dissipated in the resistor in milli-watts (mW). Round your answer to three significant figures?

4990 mW?
Answer #1

The relation between power , current and resistance is

P = I^2 R

= ( 12.2* 10^-3)^2 (33.5)

= 4.986* 10^-3 W

rounding to three significant figure

P = 4.99 mW